# RoSIS: Robust Framework for Text-Promptable Surgical Instrument Segmentation Using Vision-Language Fusion

### v1

@@ 3.3 Architecture
We design our Robust Surgical Instrument Segmentation (RoSIS) using an encoder-decoder structure to achieve robust text-promptable surgical instrument segmentation. Inspired by [15] and [14], we add a binary classifier in parallel with the decoder to determine the existence of instruments. First, we encode the image $I$ and text description $T$ using an image encoder and a text encoder, respectively. The image encoder is based on the Swin Transformer [30], which is specialized for dense prediction tasks, while the text encoder uses BERT [29]. We denote the visual feature from stage $i$ of the Swin Transformer as $v_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ and the language feature from BERT as $l\in\mathbb{R}^{T\times C_{l}}$, where $H_{i}$ and $W_{i}$ are the height and width of the visual feature, $T$ is the number of words in the text description, and $C_{i}$ and $C_{l}$ represent the number of channels for the visual and language features, respectively.
@@ 3.3.1 Encoder design
We introduce early feature fusion to efficiently integrate vision and language knowledge from the encoder. We aim to enhance semantic understanding and improve feature interactivity through early fusion. For this purpose, we add the Multi-Modal Fusion Block (MMFB) and the Selective Gate Block (SGB) between each stage of the Swin Transformer.
Fig. 3(b) shows the architecture of the MMFB, which is designed to integrate visual and language features through a multi-step fusion process, with the fused feature then passed to a gating module. The fusion process begins by independently projecting the visual and language features into compatible feature spaces. The MMFB contains two Multi-Head Cross Attention (MHCA) modules. In the first MHCA, we adopt language tokens to fuse with the language feature. Unlike [15], since we do not use the language features generated during fusion in the decoder, we focus on selecting only the information necessary for fusion with visual features. Therefore, we refine the language features with one MHCA and subsequently use them in the second MHCA to interact with the visual features.
@@ 3.3.2 Decoder design
Our decoder is designed to perform two tasks: classify the existence of the object specified by the text prompt and generate the corresponding segmentation mask. To accomplish this, we use a multi-scale deformable attention pixel decoder [31]. This decoder incorporates an FPN [32] structure to effectively utilize multi-scale features, taking $f_{1},f_{2},f_{3},$ and $f_{4}$ as inputs. The final segmentation mask $M_{c}$ is produced to pass the output from the decoder through a $1\times 1$ convolution layer.
In addition to mask generation, we introduce a parallel branch to compute the existence probability of the object described by the text prompt. This branch fuses the decoder output from $f_{4}$ with the raw language feature extracted from BERT using MHCA. This fusion enables the decoder to leverage both visual and textual information, enhancing its ability to determine the presence of the specified object. Unlike [15], which uses fused language features from the encoder for existence prediction, we use the raw language feature directly from BERT, as it retains richer semantic context. After the fusion through MHCA, the existence probability $p_{c}$ is computed by passing the output through a linear layer.
@@
Table 1: Comparison between our RoSIS with the previous vision-based and promptable method on the EndoVis2018 dataset. * means re-implemented results. $\dagger$ means that does not utilize category existence information in the image during inference.
Domain Method Robustness Ch IoU ISI IoU mc IoU Instrument category
@@
Table 2: Comparison between our RoSIS with the previous vision-based and promptable method on the EndoVis2017 dataset. * means re-implemented results. $\dagger$ means that does not utilize category existence information in the image during inference.
Domain Method Robustness Ch IoU ISI IoU mc IoU Instrument category
@@
3.4 Iterative inference
We propose an iterative inference strategy to use location prompts during inference, similar to those used in training. In the training phase, we can generate location prompts because we have access to the ground truth for the instrument, but this is not possible in the inference phase. Therefore, as shown in Fig. 4, we divide the inference process into two iterations.
In the first iteration, we generate the initial segmentation map and existence probability for each instrument category. We use two prompts: a category name prompt (e.g., “The monopolar curved scissors”) and a descriptive prompt generated by GPT, which provides visual characteristics of the instrument. For category $c$, the model predicts the segmentation maps $M_{c}^{1}\text{ and }M_{c}^{2}$ and the existence probabilities $p_{c}^{1}\text{ and }p_{c}^{2}$. If the average existence probability $(p_{c}^{1}+p_{c}^{2})/2$ exceeds 0.5, indicating that class $c$ exists in the image, the process proceeds to the second iteration.
In the second iteration, spatial information is used to refine the segmentation. Based on the segmentation maps $M_{c}^{1}\text{ and }M_{c}^{2}$ from the first iteration, the model calculates the center of mass for the detected object, allowing it to determine a rough location within one of four quadrants: left-top, left-bottom, right-top, or right-bottom. This positional information is then incorporated into a new location prompt (e.g., “The monopolar curved scissors on the right bottom”), which serves as input for the second inference. The model generates the segmentation map $M_{c}^{3}$ and the existence probability $p_{c}^{3}$ for the class $c$.
This combination leverages both the fixed prompts and the location prompt, resulting in a more precise and contextually aware segmentation map. This iterative inference approach ensures robust segmentation by progressively refining predictions based on both textual and spatial prompts, enhancing segmentation performance in complex surgical images. We describe the effects of iterative inference in detail in the ablation study in Table 5.
4 Experiment
4.1 Dataset and metric
4.2 Implementation detail
We implement our model using PyTorch [33]. For the language model, we utilize a pre-trained BERT [29]. The image encoder is initialized with an ImageNet [34] pre-trained Swin Transformer [30] to support dense prediction. In our fusion module, we set the language query size to 20. The model is trained using the AdamW [35] optimizer with a weight decay of 0.01 for 50 epochs, applying a learning rate decay of 0.1 at epochs 30 and 40. The input images, originally sized at 1280$\times$1024, are resized to 480$\times$480 for training, and the generated masks are upsampled to their original size to produce the final mask. We train the model with a batch size of 32, using four Nvidia RTX A6000 GPUs.
4.3 Results
@@ 4.3.1 Comparison on the EndoVis2018
In Table 1, we compare our method, RoSIS, with both vision-based and promptable models on the EndoVis2018 [19] dataset. For a fair comparison, we re-implement the promptable models LAVT [12] and TP-SIS [10], both of which are based on referring image segmentation tasks. For promptable methods, we provide referring sentences for all classes in the EndoVis2018 dataset, where the number of classes is fixed to seven. Pixel-wise probabilities of the generated masks were then compared to produce the final output. To ensure robustness, the vision-based models were restricted to single-image methods instead of based on the video input methods. However, S3Net is not robust because it measures the performance by selecting the top 5 instances out of 7 instances during inference. Therefore, in Table 1 and 2, we add the ’Robustness’ column to indicate that S3Net has the advantage in the inference stage.
As shown in Table 1, our method, RoSIS, not only outperforms all promptable models but also achieves superior performance compared to vision-based methods across the Ch IoU, ISI IoU, and mc IoU metrics. Vision-based models, while effective in standard segmentation tasks, do not leverage text prompts and therefore lack the adaptability provided by RoSIS’s robust promptable framework. Our method performs better even though S3Net has an advantage by limiting the number of instances during inference, with a performance improvement of +2.44% in Ch IoU and +2.89% in ISI IoU and 4.60% in mc IoU. Promptable models, such as LAVT and TP-SIS, also tend to generate a high number of false negatives, as they produce masks without first determining whether the instrument is present in the image. This limitation results in a significant gap between Ch IoU and ISI IoU scores. In contrast, our model demonstrates more stable performance by first assessing the existence of each object, effectively reducing false negatives and achieving higher segmentation accuracy.
Table 3: Text prompt design and iterative inference ablation study in EndoVis2018 dataset.
ISI IoU mc IoU
Name prompt 67.74 37.84
(+) location prompt 69.24 (+1.50) 37.54 (-0.30)
      (+) Iterative inference 73.51 (+5.77) 39.30 (+1.46)
(+) gpt4 prompt 71.37 (+3.63) 40.61 (+2.77)
      (+) Iterative inference 76.16 (+8.42) 44.54 (+6.70)
Figure 5: Visual comparison of iterative inference strategies
@@ 4.3.2 Comparison on the EndoVis2017
Table 2 presents a comparison between our method, RoSIS, and previous vision-based and promptable methods on the EndoVis2017 [18] dataset. Among the vision-based methods, S3Net [8] achieves the highest performance, surpassing our model on this dataset. However, S3Net’s inference stage is limited by a selective approach in which it only considers the top 5 instances, creating an unfair comparison and potentially omitting relevant objects. This restriction may boost its performance but reduces its applicability in real-world scenarios, where all relevant instruments need to be segmented.
@@
In contrast, our model, RoSIS, achieves competitive performance compared to S3Net with Ch IoU, ISI IoU, and mc IoU scores while operating under robust conditions. Compared to S3Net, RoSIS is designed with an end-to-end training pipeline and a straightforward inference setting. This streamlined approach makes RoSIS more suitable for practical applications, maintaining high segmentation accuracy without limiting the number of instances. Also, we evaluate other vision-based and promptable methods with RoSIS. Our method outperforms previous methods under robust conditions, has a more even overall class-wise performance distribution, and has less performance degradation for ISI IoU than for Ch IoU. This demonstrates the effectiveness of our method’s fusion structure, prompt design, and iterative inference. Notably, despite the limitations imposed by a robust evaluation environment, RoSIS achieves superior performance in two instrument categories: VS with 45.32 and UP with 44.99. These results show RoSIS’s capability to balance promptable and vision-based features, making it a versatile and efficient solution for surgical instrument segmentation.

### v3

@@ 3.3 Architecture
We design our Robust Surgical Instrument Segmentation (RoSIS) using an encoder-decoder structure to enable R-SIS. Inspired by [16] and [23], we add a binary classifier in parallel with the decoder to determine the existence of instruments. First, we encode the image $I$ and text description $T$ using an image encoder and a text encoder, respectively. The image encoder is based on the Swin Transformer [27], which is specialized for dense prediction tasks, while the text encoder uses BERT [22]. We denote the visual feature from stage $i$ of the Swin Transformer as $v_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ and the language feature from BERT as $l\in\mathbb{R}^{T\times C_{l}}$, where $H_{i}$ and $W_{i}$ are the height and width of the visual feature, $T$ is the number of words in the text description, and $C_{i}$ and $C_{l}$ represent the number of channels for the visual and language features, respectively.
@@ 3.3.1 Encoder design
We introduce early feature fusion to efficiently integrate vision and language knowledge from the encoder. We aim to enhance semantic understanding and improve feature interactivity through early fusion. For this purpose, we add the MMFB and the SGB between each stage of the Swin Transformer.
Fig. 2(b) shows the architecture of the MMFB, which is designed to integrate visual and language features through a multi-step fusion process, with the fused feature then passed to a gating module. The fusion process begins by independently projecting the visual and language features into compatible feature spaces. The MMFB contains two Multi-Head Cross Attention (MHCA) modules. In the first MHCA, we adopt language tokens to fuse with the language feature. The language token is a learnable parameter that adjusts the language feature before the fusion. Unlike [16], since we do not use the language features generated during fusion in the decoder, we focus on selecting only the information necessary for fusion with visual features. Therefore, we refine the language features with the first MHCA and subsequently use them in the second MHCA to interact with the visual features.
@@ 3.3.2 Decoder design
Our decoder is designed to perform two tasks: classify the existence of the instrument specified by the text prompt and generate the corresponding segmentation mask. To accomplish this, we use a multi-scale deformable attention pixel decoder [34]. This decoder incorporates an Feature Pyramid Networks [33] structure to effectively utilize multi-scale features, taking $f_{1},f_{2},f_{3},$ and $f_{4}$ as inputs. The final segmentation mask $M_{c}$ is produced to pass the output from the decoder through a $1\times 1$ convolution layer.
In addition to mask generation, we introduce a parallel branch to compute the existence probability of the instrument described by the text prompt. This branch fuses the decoder output from $f_{4}$ with the raw language feature extracted from BERT using MHCA. This fusion enables the decoder to leverage both visual and textual information, enhancing its ability to determine the presence of the specified instrument. Unlike [16], which uses fused language features from the encoder for existence prediction, we use raw BERT embeddings as input language features to preserve rich semantic context without introducing task-specific projection layers. While prior work adopts cross-modal embeddings aligned during training, such methods often entangle vision and language representations, which may bias prompt behavior in the presence of instrument absence. After the fusion through MHCA, the existence probability $p_{c}$ is computed by passing the output through a linear layer.
@@
Figure 3: (a) Iterative refinement process of class $c$ (monopolar curved scissors in this image): (Iteration 1) Initial segmentation using a general prompt, (Iteration 2) refined segmentation with a location prompt referring to the first predicted map. Then, we combine maps by equation 4 for accurate final output. (b) After applying the iterative refinement process to all classes, the refined maps are combined to generate the final output.
3.4 Iterative refinement for inference
We propose an iterative refinement strategy to use location prompts during inference. During training, given the ground-truth instruments present in the scene, RoSIS learns to predict segmentation masks by first determining the presence of instruments and utilizing diverse prompts, thereby reducing false positives under robust settings. During inference, where instrument existence information is not available, the model performs a two-iteration inference process, as illustrated in Fig. 3: it first predicts which instruments are visible by applying all instrument prompts, and then generates refined masks only for those instruments that are visually present in the scene.
In the first iteration, we generate the initial segmentation map and existence probability for each instrument category. We use two prompts: a category name prompt (e.g., “The monopolar curved scissors”) and a descriptive prompt generated by GPT-4, which provides visual characteristics of the instrument. For category $c$, the model predicts the segmentation maps $M_{c}^{1}\text{ and }M_{c}^{2}$ and the existence probabilities $p_{c}^{1}\text{ and }p_{c}^{2}$. If the average existence probability $(p_{c}^{1}+p_{c}^{2})/2$ exceeds 0.5, indicating that class $c$ exists in the image, the process proceeds to the second iteration.
In the second iteration, while location information can be directly obtained from the ground truth during training, it is not available during inference. Therefore, the model instead leverages the predicted masks from the first iteration to guide refinement. Based on the segmentation maps $M_{c}^{1}\text{ and }M_{c}^{2}$ from the first iteration, the model calculates the center of mass for the detected instrument, allowing it to determine a rough location within one of four quadrants: left-top, left-bottom, right-top, or right-bottom. This positional information is then incorporated into a new location prompt (e.g., “The monopolar curved scissors on the right bottom”), which serves as input for the second inference. The model generates the segmentation map $M_{c}^{3}$ and the existence probability $p_{c}^{3}$ for the class $c$.
This combination leverages both the fixed prompts and the location prompt, resulting in a more precise and contextually aware segmentation map. This iterative refinement approach ensures robust segmentation by progressively refining predictions based on both textual and spatial prompts, enhancing segmentation performance in complex surgical images. We describe the effects of iterative refinement in detail in the ablation study in Table 6.
Table 1: Prompt information for each class generated from GPT-4[15].
GPT-4 prompt information
Question Please describe the appearance of the {cls name} in endoscopic surgery with 18 words, change the description to a phrase with the subject, and do not use colons and “endoscopic surgery” and “minimally invasive procedure”.
Class name Answer
Bipolar forceps Bipolar forceps feature slim insulated handles and precise tips for targeted tissue coagulation and manipulation.
Prograsp forceps Prograsp forceps feature ergonomic handles and curved serrated jaws for secure and precise tissue manipulation during surgery.
Large needle drivers Large needle drivers are equipped with sturdy handles and sharp precise jaws for efficient suturing in surgical settings.
Monopolar curved scissors Monopolar curved scissors have slender ergonomic handles and sharp curved blades for precise cutting during surgical operations.
Ultrasound probe The ultrasound probe features a slim elongated design with a smooth tip for detailed imaging during surgical interventions.
Suction instrument The suction instrument boasts a narrow elongated tube with a controllable tip for precise fluid removal during surgery.
Clip applier The clip applier has a long slender shaft with a specialized tip for deploying clips securely during surgical procedures.
Vessel sealer Vessel sealer features a streamlined handle and specialized tips for precise sealing and division of blood vessels during surgery.
Grasping retractor Grasping retractor sports a long handle with claw-like tips for firmly holding tissues during surgical operations.
4 Experiment
4.1 Dataset and metric
As mentioned in Section 3.2, we use three fixed prompts during training for each class. The first prompt incorporates the class name, i.e., “The {cls name},” the second refers to the instrument’s location, i.e., “The {cls name} on the {location},” and the third prompt is generated from GPT-4. As shown in Table 1, the prompts generated by GPT-4 describe the appearance of the class based on its name. During inference, we only use the class name and the GPT-4 prompts, as the location information is not available at the start.
Table 2: Comparison of previous methods in terms of domain, input data type, and robustness evaluation.
Method Domain Input Robustness
TernausNet-11 [2] vision I clear
TraSeTR [4] vision S clear
MF-TAPNet [31] vision S clear
Dual-MF [32] vision S clear
ISINet [3] vision S clear
MATIS [5] vision S data leakage
S3Net [6] vision I use only top 5 instance
TP-SIS [10] promptable I oracle information
RoSIS (ours) promptable I clear
4.2 Implementation detail
We implement our model using PyTorch [30]. For the language model, we utilize a pre-trained BERT [22]. The image encoder is initialized with an ImageNet [29] pre-trained Swin Transformer [27] to support dense prediction. In our fusion module, we set the language token size to 20. The model is trained using the AdamW [28] optimizer with a weight decay of 0.01 for 50 epochs, applying a learning rate decay of 0.1 at epochs 30 and 40. The input images, originally sized at 1280$\times$1024, are resized to 480$\times$480 for training, and the generated masks are upsampled to their original size to produce the final mask. We train the model with a batch size of 32, using four Nvidia RTX A6000 GPUs.
4.3 Comparative results
Table 2 presents a comparison of previous methods in terms of domain, input data type, and robustness evaluation. The domain categorization differentiates between vision-based methods, which rely solely on image features, and promptable methods, which incorporate additional guidance such as textual prompts for segmentation. The input data type column distinguishes methods that operate on a single image (I) from those that leverage a sequence of images (S) to enhance segmentation performance using temporal information. The robustness column indicates whether each method has been evaluated under robust conditions. Methods such as MATIS [5], S3Net [6], and TP-SIS [10] do not meet robustness evaluation criteria due to issues like data leakage or oracle information. We exclude MATIS from the comparison methods because it conducts experiments excluding certain videos during the experiment. S3Net uses whole train/test data, but it selects the top 5 instances out of 7 instances during inference. TP-SIS is re-implemented under robust conditions and used in experiments.
In Tables 3 and 4, we compare our method, RoSIS, with both vision-based and promptable models on the EndoVis2018 [26] and EndoVis 2017 [25] datasets. For a fair comparison, we re-implement the promptable models LAVT [20] and TP-SIS [10], both of which are based on referring image segmentation tasks. For promptable methods, we provide referring sentences for all classes in the dataset. Pixel-wise probabilities of the generated masks were then compared to produce the final output.
Table 3: Comparison results on the EndoVis2018 dataset. * means re-implemented results. $\dagger$ means inference under not robustness conditions. Bold and underline indicate the best and the second best performance, respectively.
Domain Method Ch IoU ISI IoU mc IoU Instrument category
@@ 4.3 Comparative results
Table 4: Comparison results on the EndoVis2017 dataset. * means re-implemented results. $\dagger$ means inference under not robustness conditions. Bold and underline indicate the best and the second best performance, respectively.
Domain Method Ch IoU ISI IoU mc IoU Instrument category
@@ 4.3.1 Comparison on the EndoVis2018
As shown in Table 3, RoSIS outperforms all promptable segmentation models and achieves superior performance over vision-based methods in Ch IoU, ISI IoU, and mc IoU. S3Net is the strongest vision-based model in ISI IoU that utilizes a single image. However, RoSIS surpasses it under robust conditions, improving Ch IoU by +2.44%, ISI IoU by +2.89%, and mc IoU by +4.60%. Meanwhile, TraSeTR, which achieves the highest mc IoU among vision-based models, benefits from using sequential images instead of single-image inference. Despite this advantage, RoSIS remains competitive while relying solely on single-image processing.
Compared to promptable models like LAVT and TP-SIS, RoSIS effectively reduces false positives and false negatives. Existing promptable methods generate segmentation masks without first verifying instrument existence, causing a noticeable gap between Ch IoU and ISI IoU scores. In contrast, RoSIS incorporates an existence verification step, stabilizing the segmentation process by minimizing false negatives. This refinement results in higher segmentation accuracy, making RoSIS a more reliable solution for robust surgical instrument segmentation.
@@ 4.3.2 Comparison on the EndoVis2017
Table 4 presents a detailed comparison of segmentation performance on the EndoVis2017 [25] dataset. Among the vision-based models, S3Net [6] achieves the highest overall performance. However, this performance is achieved under a setting that retains only the top five predicted instances per frame, which may simplify the segmentation task by limiting the number of considered categories. While this assumption aligns with common surgical scenes where fewer than five instruments are present, it reduces flexibility when applied to scenarios involving a broader set of instruments or multi-phase procedures.
@@
Moreover, RoSIS outperforms other vision-based and promptable models while maintaining a more balanced class-wise performance distribution. Notably, its ISI IoU shows less performance degradation compared to Ch IoU, highlighting the effectiveness of its fusion structure, prompt design, and iterative refinement strategy. Under robust conditions, except for S3Net, RoSIS shows competitive performance overall and achieves the best performance in VS and UP. These results further validate RoSIS’s ability to effectively integrate vision-based and promptable features, making it a versatile and efficient solution for surgical instrument segmentation.
Table 5: Comparison of promptable segmentation models under the R-SIS setting using FPR, Precision, Recall, and F1-score. Bold and underline indicate the best and the second best performance, respectively.
Dataset Method FPR ↓ P ↑ R ↑ F1 ↑
EV18 LAVT [20] 0.6155 0.3770 0.3178 0.3364
RefSegformer [16] 0.2688 0.6820 0.6369 0.6420
TP-SIS [10] 0.4008 0.5374 0.4866 0.4957
RoSIS w/o exist 0.1804 0.7374 0.6714 0.6847
RoSIS (Ours) 0.0892 0.7926 0.7546 0.7611
EV17 LAVT [20] 0.7744 0.2153 0.1351 0.1483
RefSegformer [16] 0.4706 0.3505 0.2578 0.2732
TP-SIS [10] 0.3338 0.4802 0.3942 0.3981
RoSIS w/o exist 0.3532 0.5445 0.4427 0.4596
RoSIS (Ours) 0.2343 0.5907 0.5036 0.5269
4.3.3 Robustness-Oriented Comparison
Table 5 presents a comparative analysis of segmentation models under the R-SIS setting on the EndoVis2017 and EndoVis2018 datasets. We compare the performance of segmentation models under the R-SIS setting using false positive rate (FPR), precision (P), recall (R), and F1 score (F1). The results show that RoSIS substantially reduces the occurrence of false-positive predictions compared to prior promptable segmentation models. While existing methods often produce masks for absent instruments due to the assumption of instrument presence in prompts, RoSIS effectively suppresses these errors by incorporating an existence prediction mechanism. In addition to its low FPR, RoSIS achieves consistently high precision and recall, demonstrating that its robustness does not come at the cost of under-segmentation. These findings confirm the advantage of the R-SIS framework and support the use of prompt-agnostic prompting strategies for fair and reliable evaluation. Table 5 also introduces results for RoSIS without the existence prediction module (w/o exist), allowing direct comparison the effectiveness of this component. The existence prediction module leads to a clear reduction in false positives and a noticeable gain in precision across both datasets, confirming its contribution to segmentation robustness. Notably, even without the existence prediction module, RoSIS outperforms prior promptable segmentation models, indicating that its core design, rooted in robust prompting strategies and R-SIS based task framing, is inherently better suited to the evaluation conditions.
Figure 4: Visual comparison of iterative refinement strategies
Table 6: Text prompt design and iterative refinement ablation study in EndoVis2018 dataset. The values in brackets represent the increase compared to the baseline.
ISI IoU Ch IoU mc IoU
Name prompt 67.74 76.20 37.84
(+) location prompt 69.24 76.57 37.54
     (+) iterative refinement 73.51 77.58 39.30
(+) GPT-4 prompt 71.37 77.21 40.61
     (+) iterative refinement 76.16 77.66 44.54
4.4 Ablation study
4.4.1 Iterative refinement process and prompt design
Table 6 presents the ablation study results on the EndoVis2018 dataset, emphasizing the role of iterative refinement and text prompts. Starting with the baseline using only the name prompt such as “The bipolar forceps,” the model achieves ISI IoU and mc IoU scores of 67.74 and 37.84, respectively. Adding the location prompt to the name prompt slightly improves performance in ISI IoU. Then, the iterative refinement consistently improves performance across all metrics. This demonstrates that iterative refinement effectively refines the segmentation map by utilizing location-specific prompts, contributing to more accurate segmentation. Similarly, when iterative refinement is incorporated with the GPT-4 prompt, which provides descriptive characteristics for each instrument, there is a substantial improvement in segmentation accuracy. The ISI IoU increases by 4.79, and mc IoU increases by 3.93. This improvement shows the synergy between descriptive prompts and iterative refinement, as the iterative process allows the model to refine segmentation based on more detailed instrument information. Also, as shown in the red circle part of Fig. 4, iterative refinement allows for more accurate segmentation maps.
These results indicate the crucial role of iterative refinement in our model. By refining segmentation through multiple stages and adapting prompts based on previous outputs, iterative refinement contributes to significant gains in both general and class-specific accuracy, making it a key factor in achieving optimal segmentation performance.